OpenStudy (australopithecus):

A roast turkey is taken from an oven when its temperature has reached 185F and is placed on a table in a room where the temperature is 75F a) If the temperature of the turkey is 150F after half an hour what is the temperature after 45min b) when will the turkey have cooled to 100F So I got the following equation using Newtons law of cooling: 110e^((ln(75/100)/30)x) + 75 = T(t) where T(t) = the temperature of the turkey at time T For a) I got 146 Degrees F b) I got 154 minutes The text book states that the answers are a) 137 b) 116min what am I doing wrong?

OpenStudy (australopithecus):

Just rough work I set 185 = T(0) to solve for c and 150 = T(30) to solve for k

OpenStudy (mimi_x3):

\(T=A+Be^{kt}\) \(T=185, A =75, t=0\) => \(185=75+B^{0*k}\) =>\( B= 110 \) Therefore, \(T=75+110e^{kt}\) Then.. \(T=150, t=30\) \(150=75+110e^{30k}\) solve for \(k\) => \(k= -0.01\) Then the equation is \(T= 75+1110e^{-0.1k}\)

OpenStudy (mimi_x3):

typo... \(T=75+110e^{-0.1k}\)

OpenStudy (mimi_x3):

wait..another one lol \(T=75+110e^{-0.1t}\)

OpenStudy (mimi_x3):

For the temperature after \(45min\) is when you sub \(45\) into the equation and you will get the answer just like your textbook!

OpenStudy (mimi_x3):

lol, another typo my bad \(T=75+110e^{-0.01t}\)

OpenStudy (mimi_x3):

Do you understand?

OpenStudy (australopithecus):

lol

OpenStudy (australopithecus):

Didnt i get the same answer though, ln(75/100)/30) = roughly -0.001

OpenStudy (australopithecus):

sorry typo -0.01

OpenStudy (mimi_x3):

i don't know i don't undestand your solution..but you said that you didnt get teh same answer as your textbook

OpenStudy (australopithecus):

https://www.wolframalpha.com/input/?i=110e^%28%28ln%2875%2F100%29%2F30%29x%29+%2B+75 here it is in latex

OpenStudy (mimi_x3):

\[\large 110e^{\frac{1}{30}\log\left(\frac{75}{100}\right)k} +75\]?

OpenStudy (australopithecus):

yes except that k should be a t :)

OpenStudy (mimi_x3):

\[\large T= 110e^{\frac{1}{30}\log\left(\frac{75}{100}\right)t} +75\]

OpenStudy (mimi_x3):

looks like the same as mine

OpenStudy (australopithecus):

yeah can you try plugging in the numbers to see if you get the same answer

OpenStudy (australopithecus):

I think the text book is wrong but I wanted to make sure because this happens a bunch now

OpenStudy (mimi_x3):

another typo lol \[\large T= 110e^{\frac{1}{30}\log\left(\frac{75}{110}\right)t} +75\]

OpenStudy (mimi_x3):

I got the same as your textbook I assumed they round it off I got \(136.9292118\) and when you round it off its \(137\)

OpenStudy (mimi_x3):

you made a mistake its \[\huge T= 110e^{\frac{1}{30}\log\left(\frac{75}{110}\right)t} +75\]

OpenStudy (mimi_x3):

Its not \[\huge\color{ red} {T= 110e^{\frac{1}{30}\log\left(\frac{75}{100}\right)t} +75}\]

OpenStudy (australopithecus):

lol figures I made a typo my brain is garbage sometimes

OpenStudy (australopithecus):

thanks for the help mimi

OpenStudy (mimi_x3):

np